OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 74 and the last term is –100? –468 –390 –416 –442

OpenStudy (anonymous):

Sn=n(a1+an)/2=30(74-100)/2=-390

OpenStudy (lilg132):

-390